# What is the Relationship between Tensor Factorizations and Circuits (and How Can We Exploit it)?

### v1

@@ 4.1 Building and Learning Region Graphs
The first step of our pipeline is to construct a RG (Def. 4). It specifies a hierarchical partitioning of the input variables according to which we build deep circuit architectures. In particular, PCs that are built out of RGs satisfying crucial structural properties such as smoothness and decomposability by design (and structured-decomposability if the RG is a tree and has univariate leaves, see Section 2.2), which in turn guarantee tractable inference for many queries of interest (Section 2). RGs are explicitly used to build PCs in some papers (Peharz et al., 2020c; a), but as we show next, they can be implicitly found in many other PC and tensor factorization architectures. We also introduce a novel way to quickly build RGs for images that are dataset-agnostic but exploit of the structure of pixels.
@@
A simple way to instantiate a RG is by building partitionings that factorize one variable at a time. That is, given $\pi$ a ordering over variables $\bm{\mathrm{X}}$, each $i$-th partition node factorizes its scope $\{X_{\pi(1)},\ldots,X_{\pi(i)}\}$ into regions $\{X_{\pi(1)},\ldots,X_{\pi(i-1)}\}$ and $\{X_{\pi(i)}\}$. We call the resulting RG a linear tree (LT) RG, and show an example for it in case of three variables in Fig. 4. The ordering of the variables can be the lexicographical one or depending on additional information such as time when modeling sequence data. This sequential RG is the one adopted by chain-like tensor network factorizations, such as MPS, TTs or BMs (Pérez-García et al., 2007; Oseledets, 2011), as well as hidden Markov models (HMMs) when represented as PCs (Rabiner & Juang, 1986; Liu et al., 2023a).
@@ 4.2 Overparameterize & Tensorize Circuits
The first step of Algorithm 1 consists of associating input units to leaf regions, i.e., regions that are not further decomposed. Leaf regions are often univariate, i.e., of the form $\bm{\mathrm{Y}}=\{X_{j}\}$ for some variable $X_{j}\in\bm{\mathrm{X}}$. For each leaf region over a variable $X_{j}$ we introduce $K$ input units, each computing a function $f_{i}\colon\mathsf{dom}(X_{j})\to\mathbb{R}$. To guarantee the non-negativity of the output in monotonic PCs, $f_{i}$ are often chosen to be non-negative, e.g., by choosing them to be probability mass or density functions (Choi et al., 2020). However, one can possibly choose $f_{i}$ from a much wider set of expressive function families, e.g., polynomial splines (de Boor, 1971; Loconte et al., 2024a), neural networks (Shao et al., 2020; Correia et al., 2023; Gala et al., 2024a; b) and normalizing flows (Sidheekh et al., 2023). See also Section 3.2. Then, the input units can be tensorized by effectively replacing them with an input layer $\bm{\mathrm{\ell}}\colon\mathsf{dom}(X_{j})\to\mathbb{R}^{K}$ such that $\bm{\mathrm{\ell}}(X_{j})_{i}=f_{i}(X_{j})$ with $i\in[K]$ can be computed in parallel (L11 in Algorithm 1). Next, sum and product layers are built and connected according to the variables partitioning specified in the given RG.
@@ 4.3 Abstracting sum and product layers into modules
We start by considering composite layers that adopt the connectivity of computational units in the Tucker factorization, as showed in Fig. 2. This is a pattern introduced in architectures such as RAT-SPNs (Peharz et al., 2020c) and EiNets (Peharz et al., 2020a). There, a region node over $\bm{\mathrm{Y}}{\subseteq}\bm{\mathrm{X}}$ and partitioned into $(\bm{\mathrm{Z}}_{1},\bm{\mathrm{Z}}_{2})$ is parameterized as a layer $\bm{\mathrm{\ell}}$ that is a composition of a Kronecker product layer followed by a sum layer, i.e., computing
$\bm{\mathrm{\ell}}(\bm{\mathrm{Y}})=\bm{\mathrm{W}}\left(\bm{\mathrm{\ell}}_{1}(\bm{\mathrm{Z}}_{1})\otimes\bm{\mathrm{\ell}}_{2}(\bm{\mathrm{Z}}_{2})\right)$ (Tucker-layer)
@@ 4.4 Folding to Further Accelerate Learning and Inference
The final and optional step of our proposed pipeline (Fig. 8) consists of stacking together the layers that share the same functional form as to increase GPU parallelism. We name this step folding. Note that folding is only a syntactic transformation of the circuit, i.e., it does not change the encoded function and hence it preserves its expressiveness. This simple syntactic “rewriting” of a circuit can however significantly impact learning and inference performance. In fact, folding is the core ingredient of the additional speed-up introduced by EiNets (Peharz et al., 2020a) with respect to the same non-folded circuit architectures such as RAT-SPNs (Peharz et al., 2020c) which share with EiNets the other architecture details, e.g., the use of Tucker layers (see Table 1).
As such, the difference in performance that is usually reported when treating RAT-SPNs and EiNets as two different PC model classes (see e.g., Liu et al. (2023a)) must depend on other factors, such as the choice of the RG or a discrepancy in other hyperparameters used to learn these models, e.g., the chosen optimizer. By disentangling these aspects in our pipeline, we can design experiments that truly highlight which factors are responsible for increased performance (see Section 7).
@@
where $\bm{\mathrm{\ell}}_{1}$ (resp. $\bm{\mathrm{\ell}}_{2}$) denotes a folded layer computing the $F$ left (resp. right) inputs to $\bm{\mathrm{\ell}}^{(n)}$, each defined over variables $\bm{\mathrm{Z}}_{1}^{(n)}$ (resp. $\bm{\mathrm{Z}}_{2}^{(n)}$), and each $\bm{\mathrm{W}}_{n::}\in\mathbb{R}^{K\times K^{2}}$ is the parameter matrix of $\bm{\mathrm{\ell}}^{(n)}$. In other words, $\bm{\mathrm{W}}_{n::}$ is the $n$-th slice along the first dimension of a tensor $\bm{\mathcal{W}}\in\mathbb{R}^{F\times K\times K^{2}}$ obtained by stacking together the parameter matrices of each Tucker layer. Since the same region node can possibly take part in multiple partitionings of other region nodes (e.g., see Fig. 8i), we might have folded inputs $\bm{\mathrm{\ell}}_{1},\bm{\mathrm{\ell}}_{2}$ computing the same outputs. We illustrate an example of this in Fig. 8iii, which shows the folding of two Tucker sum-product layers sharing one input. For this reason, while folding provides considerable speed-ups when evaluating a tensorized circuit, it might come at the cost of increased memory usage depending on the chosen RG.
@@
It remains to decide how to choose the layers to fold together. The simplest way is traversing the tensorized circuit top-down (i.e., from the outputs towards the inputs) and to fold layers located at the same depth in the computational graph. However, note that we can also fold layers at different depth. For example, all input layers can be folded together if they encode the same input functional for all variables. This is the approach adopted in EiNets (Peharz et al., 2020a) and the one that will be used in all our experiments and benchmarks (see Section 7). However, note that this is not regarded as the optimal way to fold layers, and different ways of choosing the layers to fold might bring additional speed-ups and memory savings when tailored for specific architectures. While we do not investigate different ways of folding layers other than the one mentioned above, the disentanglement of the folding and overparameterization steps (Section 4.2) in our proposed pipeline will foster future work to rely on the wide literature on parallelizing generic computational graphs (Shah et al., 2023).
@@ 5.1 Compressing Tucker layers
where $\bm{\mathrm{\ell}}_{1},\bm{\mathrm{\ell}}_{2}$ are input layers to $\bm{\mathrm{\ell}}$ having width $K$ and scopes $\bm{\mathrm{Z}}_{1},\bm{\mathrm{Z}}_{2}$, respectively. Therefore, evaluating a compressed Tucker layer that has undergone the CP factorization requires time $\mathcal{O}(KR)$ (Eq. 18), rather than $\mathcal{O}(K^{3})$ (Eq. Tucker-layer). On top of this, we observe that if we use a CP factorization to all Tucker layers in a PC, we will obtain a circuit in which sum and product layers are not alternated anymore. For example, starting from the Tucker layers in Fig. 12(a), we would obtain a new architecture where product layers can be followed by two sum layers, as in Fig. 12(b), e.g., one parameterized by $\bm{\mathrm{A}}^{(1)}\in\mathbb{R}^{K\times R}$ feeding another sum layer parameterized by $\bm{\mathrm{B}}^{\top}\in\mathbb{R}^{R\times K}$. As we can always rewrite any composition of consecutive sum layers with a single sum layer parameterized by a product of matrices (e.g., by $\bm{\mathrm{B}}^{\top}\bm{\mathrm{A}}^{(1)}\in\mathbb{R}^{R\times R}$), we can collapse the adjacent sum layers as to obtain the simplified architecture in Fig. 12(c). More formally, under such observation and by assuming that $\bm{\mathrm{\ell}}_{1},\bm{\mathrm{\ell}}_{2}$ are also Tucker layers being decomposed, we can rewrite Eq. 18 as
@@
where $\bm{\mathrm{Q}}^{(1)},\bm{\mathrm{Q}}^{(2)}\in\mathbb{R}^{R\times R}$ are parameter matrices of sum layers, such that $\bm{\mathrm{Q}}^{(1)}=\bm{\mathrm{B}}^{\top}\bm{\mathrm{A}}^{(1)}$, $\bm{\mathrm{Q}}^{(2)}=\bm{\mathrm{C}}^{\top}\bm{\mathrm{A}}^{(2)}$.
Figure 13: CP layer.
That is, we reduced the overall width of each layer from $K$ to the smaller $R$ while still approximately computing a Tucker layer, by assuming that $\bm{\mathcal{W}}$ was originally low-rank. From now on, we will refer to Eq. CP-layer as CP layer. This is a new compositional abstraction we can use instead of Tucker layers in Algorithm 2 to build tensorized circuits out of a RG. For monotonic PCs, one can still recover the Tucker layer factorization above by replacing the CP factorization (Eq. 17) with its non-negative version (Cichocki & Phan, 2009), which ensures the factors $\bm{\mathrm{A}},\bm{\mathrm{B}},\bm{\mathrm{C}}$ and hence $\bm{\mathrm{Q}}^{(1)},\bm{\mathrm{Q}}^{(2)}$ to be non-negative matrices. Furthermore, a folded version of Eq. CP-layer can be obtained similarly to the one for Eq. Tucker-layer (see Section 4.4).
@@ 5.2 Parameter Sharing by Tensor Factorizations
We now focus on the problem of sharing parameters across layers in a tensorized PC. Again, we will exploit tensor factorizations for this task. Consider a tensorized PC built out of a RG as per our pipeline (Section 4.2). It is reasonable to assume that layers located at the same depth of the RG could store similar structure in their parameter tensors. For example, two distinct layers having adjacent pixel patches of the same size as scope may apply a similar transformation to their respective inputs, because we can assume the distributions of the two pixel patches to be quite similar. If the RG is a perfectly balanced binary tree, folding the resulting circuit translates to folding layers located at the same depth in the computational graph, i.e., layers that are likely to share the same structure in parameter space. This motivates us to implement parameter sharing as a factorization across layers being folded.
Specifically, we start by compressing a folded Tucker layer (Eq. Tucker-folded) and to retrieve a new layer that implements the aforementioned parameter sharing, we again decompose its parameter tensor via a CP factorization (Def. 10). This time, we will have to decompose $\bm{\mathcal{W}}\in\mathbb{R}^{F\times K\times K\times K}$, i.e., the fourth-order tensor obtained by reshaping of the parameter tensor of a folded Tucker layer $\bm{\mathrm{\ell}}$, where $F$ indicates the folding dimension. By applying a rank-$R$ CP factorization such that $R\ll K$, we obtain that
@@
where $\bm{\mathrm{Q}}^{(1)},\bm{\mathrm{Q}}^{(2)}\in\mathbb{R}^{R\times R}$ do not depend on the fold dimension, and $\bm{\mathrm{D}}\in\mathbb{R}^{F\times R}$. However, we can go further in sharing parameters and drop the fold-dependent parameter matrix $\bm{\mathrm{D}}$ from Eq. CPS-layer, hence effectively fixing it to be a matrix of ones. The reason is that its contribution can be “absorbed” by the matrices associated to the following sum layers (i.e., similarly to the “collapse” of consecutive sum layers shown in Fig. 12). We will refer to this layer as CPXS. Our experiments (Section 7) support this conjecture: as we experiment with both CPS and CPXS, we find they achieve comparable performances for distribution estimation. These two new composite layers are a nice addition to the circuit literature, and possible inspiration to further layer designs.
@@
Opportunity 7. Many new layer and circuit architectures So far we introduced Tucker, CP, $\text{CP}^{\top}$, CPS and CPXS as possible composite layers for circuits and tensor factorizations. However, one is not limited to this list and can design new ones: as long as they are compositions of the same building blocks outlined in Def. 7, they can be seamlessly plugged into Algorithm 1. Our experiments (Section 7) show the combination of RG and layer choices can have a significant impact on the resulting performances of the chosen architecture (may it be time and memory requirements or accuracy as distribution estimators), hence justifying further exploring the design space of PC architectures. Lastly, we remark that one is not limited to pick the same composite layer for each node in a RG, according to Algorithm 1. From the point of view of tensor factorizations, this would result in a peculiar “Frankenstein” hierarchical tensor factorization that mixes different local factorizations. From a ML perspective, determining which layer structure to select for each RG node can be cast as a neural architecture search task (Ren et al., 2021).
6 Additional Related Work
Tensor networks and PGMs.
TNs (Orús, 2013) are widely used to model many-body systems in physics and quantum mechanics (Schollwoeck, 2010), and have been used to simulate quantum computations on classical hardware (Markov & Shi, 2008). They have been applied more recently for machine learning applications (Stoudenmire & Schwab, 2016; Han et al., 2018; Efthymiou et al., 2019; Bonnevie & Schmidt, 2021). As they essentially an alternative formalism for probabilistic graphical models over discrete variables (Koller & Friedman, 2009), people have started drawing connections between the two formalisms. For example, Bonnevie & Schmidt (2021) connects non-negative MPS/TTs to PGMs and offers routines for probabilistic reasoning. Similarly, Glasser et al. (2020) explores the same connection, but instead of drawing TNs as PGMs, they draw them as factor graphs (Kschischang et al., 2001).
Instead, by representing non-negative tensor factorizations as monotonic PCs, we effortlessly unlock the developed theory and algorithms required to perform complex probabilistic inference, with possible applications in lossless compression (Liu et al., 2022), neuro-symbolic AI with correctness guarantees (Ahmed et al., 2022) and constrained text generation (Zhang et al., 2023). Finally, results about the succinctness or expressive efficiency of these factorizations (Glasser et al., 2019) have been used recently to prove circuit lowerbounds (Loconte et al., 2024a; b).
Probabilistic circuits and PGMs.
The modern formulation of PCs has been introduced for the first time in (Vergari et al., 2019b) as a unifying framework for several existing tractable probabilistic models (TPMs) including arithmetic circuits, (Darwiche, 2001), probabilistic decision graphs (Jaeger, 2004), and-or graphs (Marinescu & Dechter, 2009), cutset networks (Rahman et al., 2014), sum-product networks (Poon & Domingos, 2011) and more (Choi et al., 2020). The aim of PCs has been to abstract away from the different syntaxes and model formalisms of the above TPMs and focus on structural properties that enable tractable inference in each. Non-negative tensor factorizations and tensor networks have been underlooked in this effort so far. Several ways to compile discrete PGMs into PCs (or one of the above formalisms) have been devised in the past (Oztok & Darwiche, 2017; Shen et al., 2016; Choi et al., 2013). These compilation techniques yield sparse deterministic circuits, and only recently PCs have started to be represented first in code (Peharz et al., 2020c; a; Liu & Van den Broeck, 2021b) and then formally (Loconte et al., 2024a) as tensorized architectures. Perhaps this lack of tensorized compilation targets has hidden the connection between PCs and matrix and tensor factorizations. The closest connection we are aware of can be found in Jaini et al. (2018b): they bridge sum-product networks to hierarchical mixture models and HMMs and hint at a connection with tensorial mixture models (Sharir et al., 2017) a variant of hierarchical Tucker (Def. 5). Loconte et al. (2024a) formally reduced MPSs and BMs to circuits and started drawing a bridge with the literature on expressiveness of tensor factorizations (Glasser et al., 2019).
Matrix factorizations and circuit complexity results.
Finding lower bounds to the rank of matrix factorizations can be used as a proxy to prove lower bounds to the size of circuits satisfying particular structural properties (de Colnet & Mengel, 2021). Proving an exponential (w.r.t. the number of variables) size lower bound for a class of circuits shows a limitation on which functions they can compute in polynomial time and number of parameters, thus allowing us to precisely separate circuit classes in terms of their expressiveness (Valiant, 1979; Martens & Medabalimi, 2014). Recently, lower bounding the non-negative rank (Gillis, 2020) and the square root rank (Fawzi et al., 2014; Lee & Wei, 2014) of matrices has been used to draw an expressiveness hierarchy of classes of PCs with negative real and complex-valued parameters for distribution estimation (Loconte et al., 2024a; b). Since circuits generalize many tensor network factorizations (see Section 2.4), showing size lower bounds for a class of circuits can be used to show size lower bounds for tensor networks regardless of their structure, e.g. as shown by Loconte et al. (2024b) in generalizing a known rank lower bound for real Born machines obtained by squaring a MPS/TT (Glasser et al., 2019).
7 Empirical Evaluation: Which RG and Layers to use?
@@
Table 2: Distribution estimation results. We report the test-set bpd of our best architectures, QT-CP-512 and QG-CP-512, and compare them against HCLT (Liu & Van den Broeck, 2021b), RAT-SPN (Peharz et al., 2020c), SparsePC (Dang et al., 2022a), IDF (Hoogeboom et al., 2019), BitSwap (Kingma et al., 2019), BBans (Townsend et al., 2019) and McBits (Ruan et al., 2021). SparsePC is a structure learning algorithm for PCs that iteratively finetunes both structure and parameters of a trained PC and can potentially be applied as a post-processing step to the PCs we are learning with our pipeline. HCLT results are taken from (Gala et al., 2024a). Dataset CelebA* is preprocessed using the lossless YCoCg transform.

### v2

@@ 4.1 Building and Learning Region Graphs
The first step of our pipeline is to construct a RG (Def. 4). It specifies a hierarchical partitioning of the input variables according to which we build deep circuit architectures. In particular, PCs that are built out of RGs satisfying crucial structural properties such as smoothness and decomposability by design (and structured-decomposability if the RG is a tree and has univariate leaves, see Section 2.2), which in turn guarantee tractable inference for many queries of interest (Section 2). RGs are explicitly used to build PCs in some papers (Peharz et al., 2020c; a), but as we show next, they can be implicitly found in many other PC and tensor factorization architectures. We also introduce a novel way to quickly build RGs for images that are dataset-agnostic but exploit the structure of pixels.
@@
A simple way to instantiate a RG is by building partitionings that factorize one variable at a time. That is, given an ordering $\pi$ over variables $\bm{\mathrm{X}}$, each $i$-th partition node factorizes its scope $\{X_{\pi(1)},\ldots,X_{\pi(i)}\}$ into regions $\{X_{\pi(1)},\ldots,X_{\pi(i-1)}\}$ and $\{X_{\pi(i)}\}$. We call the resulting RG a linear tree (LT) RG, and show an example for it in case of three variables in Fig. 3. The ordering of the variables can be the lexicographical one or depending on additional information such as time when modeling sequence data. This sequential RG is the one adopted by chain-like tensor network factorizations, such as MPS, TTs or BMs (Pérez-García et al., 2007; Oseledets, 2011), as well as hidden Markov models (HMMs) when represented as PCs (Rabiner & Juang, 1986; Liu et al., 2023a).
@@ 4.2 Overparameterize & Tensorize Circuits
The first step of Algorithm 1 consists of associating input units to leaf regions, i.e., regions that are not further decomposed. Leaf regions are often univariate, i.e., of the form $\bm{\mathrm{Y}}=\{X_{j}\}$ for some variable $X_{j}\in\bm{\mathrm{X}}$. For each leaf region over a variable $X_{j}$ we introduce $K$ input units, each computing a function $f_{i}\colon\mathsf{dom}(X_{j})\to\mathbb{R}$. To guarantee the non-negativity of the output in monotonic PCs, $f_{i}$ are often chosen to be non-negative, e.g., by choosing them to be probability mass or density functions (Choi et al., 2020). However, one can possibly choose $f_{i}$ from a much wider set of expressive function families, e.g., polynomial splines (de Boor, 1971; Loconte et al., 2024), neural networks (Shao et al., 2020; Correia et al., 2023; Gala et al., 2024a; b) and normalizing flows (Sidheekh et al., 2023). See also Section 3.2. Then, the input units can be tensorized by effectively replacing them with an input layer $\bm{\mathrm{\ell}}\colon\mathsf{dom}(X_{j})\to\mathbb{R}^{K}$ such that $\bm{\mathrm{\ell}}(X_{j})_{i}=f_{i}(X_{j})$ with $i\in[K]$ can be computed in parallel (L11 in Algorithm 1). Next, sum and product layers are built and connected according to the variables partitioning specified in the given RG.
@@ 4.3 Abstracting Sum and Product Layers into Modules
We start by considering composite layers that adopt the connectivity of computational units in the Tucker factorization, as shown in Fig. 2. This is a pattern introduced in architectures such as RAT-SPNs (Peharz et al., 2020c) and EiNets (Peharz et al., 2020a). There, a region node over $\bm{\mathrm{Y}}{\subseteq}\bm{\mathrm{X}}$ and partitioned into $(\bm{\mathrm{Z}}_{1},\bm{\mathrm{Z}}_{2})$ is parameterized as a layer $\bm{\mathrm{\ell}}$ that is a composition of a Kronecker product layer followed by a sum layer, i.e., computing
$\bm{\mathrm{\ell}}(\bm{\mathrm{Y}})=\bm{\mathrm{W}}\left(\bm{\mathrm{\ell}}_{1}(\bm{\mathrm{Z}}_{1})\otimes\bm{\mathrm{\ell}}_{2}(\bm{\mathrm{Z}}_{2})\right),$ (Tucker-layer)
@@ 4.4 Folding to Further Accelerate Learning and Inference
The final and optional step of our proposed pipeline (Fig. 9) consists of stacking together the layers that share the same functional form as to increase GPU parallelism. We name this step folding. Note that folding is only a syntactic transformation of the circuit, i.e., it does not change the encoded function and hence it preserves its expressiveness. This simple syntactic “rewriting” of a circuit can however significantly impact learning and inference performance. In fact, folding is the core ingredient of the additional speed-up introduced by EiNets (Peharz et al., 2020a) with respect to the same non-folded circuit architectures such as RAT-SPNs (Peharz et al., 2020c) which share with EiNets the other architecture details, e.g., the use of Tucker layers (see Table 1). As such, the difference in performance that is usually reported when treating RAT-SPNs and EiNets as two different PC model classes (see e.g., Liu et al. (2023a)) must depend on other factors, such as the choice of the RG or a discrepancy in other hyperparameters used to learn these models, e.g., the chosen optimizer. By disentangling these aspects in our pipeline, we can design experiments that truly highlight which factors are responsible for increased performance (see Section 6).
@@
where $\bm{\mathrm{\ell}}_{1}$ (resp. $\bm{\mathrm{\ell}}_{2}$) denotes a folded layer computing the $F$ left (resp. right) inputs to $\bm{\mathrm{\ell}}^{(n)}$, each defined over variables $\bm{\mathrm{Z}}_{1}^{(n)}$ (resp. $\bm{\mathrm{Z}}_{2}^{(n)}$), and each $\bm{\mathrm{W}}_{n::}\in\mathbb{R}^{K\times K^{2}}$ is the parameter matrix of $\bm{\mathrm{\ell}}^{(n)}$. In other words, $\bm{\mathrm{W}}_{n::}$ is the $n$-th slice along the first dimension of a tensor $\bm{\mathcal{W}}\in\mathbb{R}^{F\times K\times K^{2}}$ obtained by stacking together the parameter matrices of each Tucker layer. Since the same region node can possibly take part in multiple partitionings of other region nodes (e.g., see Fig. 9i), we might have folded inputs $\bm{\mathrm{\ell}}_{1},\bm{\mathrm{\ell}}_{2}$ computing the same outputs. We illustrate an example of this in Fig. 9iii, which shows the folding of two Tucker sum-product layers sharing one input. In Appendix F, we report a pytorch snippet implementing a folded Tucker layer with an einsum operation. For this reason, while folding provides considerable speed-up when evaluating a tensorized circuit, it might come at the cost of increased memory usage depending on the chosen RG.
@@
It remains to decide how to choose the layers to fold together. The simplest way is traversing the tensorized circuit top-down (i.e., from the outputs towards the inputs) and to fold layers located at the same depth in the computational graph. However, note that we can also fold layers at different depth. For example, all input layers can be folded together if they encode the same input functional for all variables. This is the approach adopted in EiNets (Peharz et al., 2020a) and the one that will be used in all our experiments and benchmarks (see Section 6). However, note that this is not regarded as the optimal way to fold layers, and different ways of choosing the layers to fold might bring additional speed-up and memory savings when tailored for specific architectures. While we do not investigate different ways of folding layers other than the one mentioned above, the disentanglement of the folding and overparameterization steps (Section 4.2) in our proposed pipeline will foster future work to rely on the wide literature on parallelizing generic computational graphs (Shah et al., 2023).
@@ 5.1 Compressing Tucker layers
where $\bm{\mathrm{\ell}}_{1},\bm{\mathrm{\ell}}_{2}$ are input layers to $\bm{\mathrm{\ell}}$ having width $K$ and scopes $\bm{\mathrm{Z}}_{1},\bm{\mathrm{Z}}_{2}$, respectively.
Figure 14: CP layer.
Therefore, evaluating a compressed Tucker layer that has undergone the CP factorization requires time $\mathcal{O}(KR)$ (Eq. 16), rather than $\mathcal{O}(K^{3})$ (Eq. Tucker-layer). On top of this, we observe that if we use a CP factorization to all Tucker layers in a PC, we will obtain a circuit in which sum and product layers are not alternated anymore. For example, starting from the Tucker layers in Fig. 13(a), we would obtain a new architecture where product layers can be followed by two sum layers, as in Fig. 13(b), e.g., one parameterized by $\bm{\mathrm{A}}^{(1)}\in\mathbb{R}^{K\times R}$ feeding another sum layer parameterized by $\bm{\mathrm{B}}^{\top}\in\mathbb{R}^{R\times K}$. As we can always rewrite any composition of consecutive sum layers with a single sum layer parameterized by a product of matrices (e.g., by $\bm{\mathrm{B}}^{\top}\bm{\mathrm{A}}^{(1)}\in\mathbb{R}^{R\times R}$), we can collapse the adjacent sum layers as to obtain the simplified architecture in Fig. 13(c). More formally, under such observation and by assuming that $\bm{\mathrm{\ell}}_{1},\bm{\mathrm{\ell}}_{2}$ are also Tucker layers being decomposed, we can rewrite Eq. 16 as
@@
where $\bm{\mathrm{Q}}^{(1)},\bm{\mathrm{Q}}^{(2)}\in\mathbb{R}^{R\times R}$ are parameter matrices of sum layers, such that $\bm{\mathrm{Q}}^{(1)}=\bm{\mathrm{B}}^{\top}\bm{\mathrm{A}}^{(1)}$, $\bm{\mathrm{Q}}^{(2)}=\bm{\mathrm{C}}^{\top}\bm{\mathrm{A}}^{(2)}$. That is, we reduced the overall width of each layer from $K$ to the smaller $R$ while still approximately computing a Tucker layer, by assuming that $\bm{\mathcal{W}}$ was originally low-rank. From now on, we will refer to Eq. CP-layer as CP layer. This is a new compositional abstraction we can use instead of Tucker layers in Algorithm 2 to build tensorized circuits out of a RG. For monotonic PCs, one can still recover the Tucker layer factorization above by replacing the CP factorization (Eq. 15) with its non-negative version (Cichocki & Phan, 2009), which ensures the factors $\bm{\mathrm{A}},\bm{\mathrm{B}},\bm{\mathrm{C}}$ and hence $\bm{\mathrm{Q}}^{(1)},\bm{\mathrm{Q}}^{(2)}$ to be non-negative matrices. Furthermore, a folded version of Eq. CP-layer can be obtained similarly to the one for Eq. Tucker-layer (see Section 4.4).
@@ 5.2 Parameter Sharing by Tensor Factorizations
We now focus on the problem of sharing parameters across layers in a tensorized PC. We again exploit tensor factorizations for this task. Consider a tensorized PC built out of a RG as per our pipeline (Section 4.2). It is reasonable to assume that layers located at the same depth might store a similar structure in their parameter tensors. For example, two distinct layers having adjacent pixel patches of the same size as scope may apply a similar transformation to their respective inputs, as we can assume the distributions of the two pixel patches to be quite similar. If the RG is a perfectly balanced binary tree, folding the resulting circuit translates to folding layers located at the same depth, which are likely to share similar structure in parameter space. This motivates us to implement parameter sharing as a factorization across folded layers.
Specifically, we start by compressing a folded Tucker layer (Eq. Tucker-folded) and to retrieve a new layer that implements the aforementioned parameter sharing, we again decompose its parameter tensor via a CP factorization (Def. 10). This time, we will have to decompose $\bm{\mathcal{W}}\in\mathbb{R}^{F\times K\times K\times K}$, i.e., the 4-dimensional tensor obtained by reshaping of the parameter tensor of a folded Tucker layer $\bm{\mathrm{\ell}}$, where $F$ indicates the folding dimension. By applying a rank-$R$ CP factorization such that $R\ll K$, we obtain that
@@
where $\bm{\mathrm{Q}}^{(1)},\bm{\mathrm{Q}}^{(2)}\in\mathbb{R}^{R\times R}$ do not depend on the fold dimension, and $\bm{\mathrm{D}}\in\mathbb{R}^{F\times R}$. However, we can go further in sharing parameters and drop the fold-dependent matrix $\bm{\mathrm{D}}$ from Eq. CPS-layer, hence effectively fixing it to be a matrix of ones. The reason is that its contribution can be “absorbed” by the matrices associated to the following sum layers (i.e., similarly to the “collapse” of consecutive sum layers shown in Fig. 13). We refer to this layer as CPXS. Our experiments (Section 6) support this conjecture: as we experiment with both CPS and CPXS, we find they achieve comparable performances for distribution estimation. These new layers are a nice addition to the circuit literature, and possible inspiration for further layer designs.
@@
Opportunity 7. Many new layers and circuit architectures We introduced Tucker, CP, $\text{CP}^{\top}$, CPS and CPXS as possible composite layers for circuits and tensor factorizations, and we provide PyTorch snippets of them in Appendix F. However, one is not limited to such layers and can design new ones: as long as they are compositions of the building blocks outlined in Def. 7, they can be seamlessly plugged into Algorithm 1 to construct new tensor factorizations as tensorized circuits. Our experiments (Section 6) show that the choice of RG and layer significantly impacts the performance of the resulting architecture (may it be time and memory requirements or accuracy as distribution estimators), hence justifying further exploration of the design space of PC architectures. Lastly, we remark that one is not limited to pick the same composite layer for each node in a RG, according to Algorithm 1. From the point of view of tensor factorizations, this would result in a peculiar “Frankenstein” hierarchical tensor factorization that mixes different local factorizations, as shown in Fig. 6. From an ML perspective, determining which layer structure to select for each RG node can be cast as a neural architecture search task (Ren et al., 2021).
@@
Table 2: Distribution estimation results. We report the test-set bpd of our best architectures, QT-CP-512 and QG-CP-512, and compare them against HCLT (Liu & Van den Broeck, 2021b), RAT-SPN (Peharz et al., 2020c), SparsePC (Dang et al., 2022a), IDF (Hoogeboom et al., 2019), BitSwap (Kingma et al., 2019), BBans (Townsend et al., 2019) and McBits (Ruan et al., 2021). SparsePC is a structure learning algorithm for PCs that iteratively finetunes both structure and parameters of a trained PC and can potentially be applied as a post-processing step to the PCs we are learning with our pipeline. HCLT results are taken from Gala et al. (2024a). Dataset CelebA* is preprocessed using the lossless YCoCg transform.